Obtain the value of the constant c for which the function u(x,t) = cosαcxsinαt is a
solution of the wave equation :
d^2u/dt^2=c^2(d^2u/dx^2)
1
Expert's answer
2019-03-15T12:38:10-0400
"d^2 u\/dt^2= - \u03b1^2 cos\u03b1cxsin\u03b1t"

"d^2u\/dx^2=-\u03b1^2c^2cos\u03b1cxsin\u03b1t"

Thus,


"-α^2=c^2(-α^2c^2)\implies c^4=1"

c=+1 or c=-1